Solve (t+10)^2-36=0, where t is a real number. Simplify your answer as much as possible, if more than 1 solution separate with a comma
Answer by bobdabuildr(27)

(t+10)^2=36 Find the sqrt of both sides
(t+10)=6 Open the brackets
t+10=6 Subtract 10 from both sides
t=-4